The air mass of 0.05 kg works on a Carnot machine. The maximum temperature of
cycle is 940 ° K and the maximum pressure is 8.4 x 10 ^ 3 Kpa. The heat supplied by
cycle is 4.2 Kj. Determine the maximum volume of the cylinder if the minimum temperature.
during the cycle it is 300 ° k
Answer #1

Is = H.PKG heat addition in 2-3 T=0 0,09kg as pl T 1. quo K PE P &nxiora رم 84 in T-7= 300* vi 2 v w 4 is minim volune da volBook pr= mgT 8TH c.os70.227x300 7 1.86H503 V = 2.25.26 m3 4 maximum volume.